A jet plane at take-off can produce sound of intensity 9.30 W/m2 at 30.4 m away. But you prefer the tranquil sound of normal conversation, which is 1.0 μW/m2. Assume that the plane behaves like a point source of sound.

What intensity from the jet does your friend experience if she lives twice as far from the runway as you do?
Answer #1

At any distance the area is proportional to the square of the radius.
So I2/I1 = (R1 / R2)^2

for the same reason you double the distance and you make the volume 1/ 2^2 = 1/4

Your friend will experience 1/4 the sound energy per m^2 ( inversely proportional to the square of the distance )

I = 0.25 μW/m2